# On the Mathematical Foundation of a Decoupled Directional Distortional Hardening Model for Metal Plasticity in the Framework of Rational Thermodynamics

### v2

@@ Abstract
This study proposes a modification to the yield condition that addresses the mathematical constraints inherent in the Directional Distortional Hardening models developed by Feigenbaum and Dafalias. The modified model resolves both the mathematical inconsistency found in the “complete model” and the limitations of the “r-model.” In the complete model, inconsistency arises between the distortional term in the yield surface and the plastic part of the free energy in the absence of kinematic hardening. Additionally, the r-model fails to capture the flattening of the yield surface in the reverse loading direction due to the absence of a fourth-order anisotropic tensor structure in the distortional term. To address these issues, the proposed model introduces a decoupled distortional hardening term in the yield function. This modification enables the simultaneous representation of both flattening and sharpening of the yield surface, and permits isotropic hardening with distortion even without kinematic hardening. A consistent mathematical formulation based on rational mechanics and a corresponding numerical algorithm are also developed, establishing a foundation for future experimental investigations and model validation.
@@
In memory of Jacob Lubliner
@@ 1 Introduction
Various attempts have been made over the years to mathematically model the yield surface to explain the phenomenon of Directional Distortional Hardening. One of the most notable works is by Baltov and Sawczuk [6] where the author introduced a fourth-order anisotropic tensor in the functional representation of yield surface. This fourth-order anisotropic tensor is the internal variable used to model the distortion. Dafalias et al. [7] proposed a yield surface model, where anisotropy is developed through kinematic and distortional hardening and distortion is represented by a fourth-order tensor $\mathbf{H}$. In this model, the evolution equation of the internal variables was developed using the principles of thermodynamics. Voyiadjis and Foroozesh [8] used the same yield surface as Dafalias et al. [7]. However, the fourth-order distortional tensor $\mathbf{H}$ is defined differently. Voyiadjis and Foroozesh [8] used kinematic considerations to develop the evolution equation for this fourth-order tensor unlike the work of Dafalias et al. [7] where thermodynamic considerations were used in this purpose. Other attempts to model directional distortion have been made by Ortiz and Popov [9], Kurtyka and Zyczkowski [10] and François [11]. Ortiz and Popov [9] used a Fourier expansion of the harmonic function to model the yield surface. Kurtyka and Zyczkowski [10] used five hyperspheres with different radius and centers to model the distortion of the yield surface. François [11] included the directional distortion term within the context of thermodynamics for the first time. Unlike the models relying on fourth-order internal variables, his approach introduced distortion by redefining the stress measure itself using a distorted stress formulation governed by second-order ranked internal variables to describe yield surface evolution under non-proportional loading. Feigenbaum and Dafalias [12] developed a ‘complete model’ based on the work of Dafalias et al. [7]. In the ‘complete model’ [12], Feigenbaum incorporated the directionality of the distortion into the yield surface equation. The equations used in this model are derived based on the second law of thermodynamics, specifically the dissipation inequality. This model suggests that energy is released during directional distortion, stored during kinematic and isotropic hardening, and uses evanescent memory type hardening rules. However, the model relies on the back-stress tensor to determine the directionality of distortion, and hence when back-stress equals zero, the model reduces to an isotropic von Mises type model, even if the anisotropy tensor $\mathbf{A}$ is non-zero. And the ‘r-model’ [13] is also developed subsequently by Feigenbaum and Dafalias [13] which introduced an orientational second-order tensor ($\mathbf{r}$) instead of the back-stress ($\boldsymbol{\alpha}$) for the scalar multiplier. The ‘r-model’ replaces the fourth-order distortion tensor $\mathbf{H}$ with the scalar quantity $(1-\mathbf{n^{r}}:\mathbf{r})$, where $\mathbf{n^{r}}$ is the unit radial tensor from the center of the yield surface to the current stress state and $(\mathbf{n^{r}}:\mathbf{r})$ is completely responsible for directional distortion. Therefore, it becomes possible to decouple kinematic hardening from distortional hardening in this model. The mathematical expression for $\mathbf{n}_{r}=\frac{\mathbf{s}-\boldsymbol{\alpha}}{\lVert\mathbf{s}-\boldsymbol{\alpha}\rVert}$, where $\mathbf{s}$ is the deviatoric part of the stress tensor and $\boldsymbol{\alpha}$ is the backstress tensor. Due to the absence of the fourth-order anisotropic term $\mathbf{A}$ in the equation of the yield surface, the flattening of the yield surface in the opposite direction of loading cannot be captured using this model.
[…]
@@ 2 Framework of Feigenbaum and Dafalias Distortional Hardening Model
$H_{0ijkl}=\frac{3}{2}I_{ijkl}=\frac{3}{2}\left[\frac{1}{2}\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}\right)-\frac{1}{3}\delta_{ij}\delta_{kl}\right].$ (3)
@@
[…]
@@ 2.1 Complete Model and Its Inconsistency
The ‘complete model’ [12] is based on Coleman Noll type rational thermodynamics [18] and includes a fourth-order tensor-valued internal variable ($\mathbf{A}$) that describes the evolving directional distortion. Hardening equations for the internal variables, as given in Eqs. (7) to (9), are derived based on sufficient conditions to satisfy the thermodynamic requirements of positive energy dissipation. Evolution equtions of the internal variables are all fading memory-type equations.
@@
The isotropic, kinematic, and distortional parts of the free energy are positive definite, and their evolution equations asymptotically approach a saturation limit. In addition, if the internal variable of a hardening term is zero, the corresponding free energy also becomes zero. For example, if $\boldsymbol{\alpha}=\boldsymbol{0}$, $\psi_{p}^{\text{kin}}=\frac{1}{2a_{1}}\boldsymbol{\alpha}:\boldsymbol{\alpha}=0$, and Eq. (12)7 reduces to Eq. (13):
@@
The term $[1-(\mathbf{n}^{r}:\mathbf{r})]$ introduces directionality to the yield surface, where,$\mathbf{n^{r}}$ is the normalized ($\boldsymbol{s-\alpha}$) in the radial direction as defined in Eq. 4 and $\mathbf{r}$ is a second order tensor-valued internal variable describing the preferred direction of distortion. As a result, the kinematic hardening has been decoupled from distortional hardening in this model.
@@
Because of this simplification, it is possible to model distortion and translation of the yield surface separately and simultaneously, which was a major issue in the complete model. The mathematical form of the yield surface in Eq. (15) is motivated by the experiments of Ishikawa [19], which suggest that the yield surface starts as a ‘von Mises’ hyper-circle, and the trace-type scalar term $(\mathbf{n^{r}}:\mathbf{r})$ distorts the yield surface in the direction of loading. However, the flattening in the opposite side of the yield surface, as observed in the experiments [1, 2, 3, 4], cannot be modeled using this ‘r-model’. A yield surface model with a fourth-order anisotropic tensor could provide an effective means to capture this flattening phenomenon.
@@ 3.2 Hardening Rules based on the Proposed Model
Based on the proposed form of the yield surface in Eq. (17), the evolution equations for this model and the relevant expressions are derived from the second law of thermodynamics as discussed below.
@@ 3.2.1 Expression of $\frac{\partial f}{\partial\boldsymbol{\sigma}}$ and Flow Rule
$\frac{\partial f}{\partial\boldsymbol{\sigma}}=||\mathbf{s}-\boldsymbol{\alpha}||\{3\mathbf{n}^{r}+2(\mathbf{n}^{r}:\mathbf{r})(\mathbf{r}\otimes\mathbf{r}):\mathbf{n}^{r}+(\mathbf{n}^{r}:(\mathbf{r}\otimes\mathbf{r})):\mathbf{n}^{r}(\mathbf{r}-(\mathbf{n}^{r}:\mathbf{r})\mathbf{n}^{r})\}.$ (18)
@@ 3.2.2 Derivation of Hardening Rules
$\psi_{p}^{\text{iso}}=\sum_{i=1}^{n}\frac{\kappa_{i}}{2\rho}k_{c}^{2},$ (21)
$\psi_{p}^{\text{kin}}=\sum_{i=1}^{n}\frac{a_{i}}{2\rho}\boldsymbol{{\alpha_{c}}}:\boldsymbol{{\alpha_{c}}},$ (22)
$\psi_{p}^{\text{dis}}=\sum_{i=1}^{n}\frac{A_{i}}{2\rho}\mathbf{r_{c}}:\mathbf{r_{c}}.$ (23)
To keep the calculation simple and the number of parameters to the minimum, only the first term is considered in Eqs. (21)–(23), i.e., $n=1$, and no superposition is considered. The final expressions are given in Eqs. (24)–(26):
@@
$\dot{\boldsymbol{\varepsilon}^{p}}-\frac{1}{a_{1}}\dot{\boldsymbol{\alpha}}=\lambda\left\|\frac{\partial f}{\partial\boldsymbol{\sigma}}\right\|a_{2}\boldsymbol{\alpha}\quad\implies\quad\dot{\boldsymbol{\alpha}}=\lambda\left\|\frac{\partial f}{\partial\boldsymbol{\sigma}}\right\|a_{1}\left(\mathbf{n}-a_{2}\boldsymbol{\alpha}\right),$ (35)
$\lambda\,||\mathbf{s}-\boldsymbol{\alpha}||^{2}\Bigl{[}(\mathbf{n^{r}}:\mathbf{r})\,\mathbf{n^{r}}:(\mathbf{r}\otimes\mathbf{r}):\mathbf{n^{r}}\Bigr{]}+\frac{1}{A_{1}}\,\mathbf{r}:\boldsymbol{\dot{r}}=-\frac{3}{2}\lambda\,||\mathbf{s}-\boldsymbol{\alpha}||^{2}\,A_{2}\,(\mathbf{r}:\mathbf{r}),\$
@@
The rate equations obtained are of the evanescent memory type, i.e., these are the Armstrong-Frederick type equations. This means the internal variables have an upper limit on their evolution and are path-dependent. For Eq. (36) to satisfy Eq. (33), the necessary condition is given in Eq. (40):
@@ 3.2.3 Plastic Modulus
$\frac{\partial f}{\partial\boldsymbol{\alpha}}=-\frac{\partial f}{\partial\boldsymbol{\sigma}}.$ (43)
@@ 3.2.4 Constraints on Material Parameters
Another way to find and cross-check the limit is by substituting Eqs. (4749) into the expression of the gradient of the yield surface at the limiting condition. Substituting the limiting values of the internal variables obtained in Eqs. (4749) into the expression of the gradient of the yield surface given in Eq. (18), the limiting value of the gradient is obtained and given in Eq. (51):
@@
Remarkably, the thermodynamic approach and the convexity requirement of the yield surface both agree upon the restriction put on the material constant. This can also be calculated from the positive definiteness of the equivalent stress which enters the yield function, $f$ at the limiting case of $\mathbf{n}^{r}$ and $\boldsymbol{\alpha}$ can also be calculated as expressed in Eq. (54):
@@
Which disagrees with Eqs. (50) and (53). Therefore, $\mathbf{n^{r}}:\mathbf{n^{r}}^{l}=+1$ must be the case, i.e., they act in the same direction. In that case, it agrees with Eqs. (50) and (53) and satisfies the inequality in Eq. (56).
@@
4 Numerical Experiment
The goal of this section is to show that the theoretical model can be implemented numerically using the algorithm described in the appendix. To show this the material is loaded uniaxially with a strain increment of $5\times 10^{-5}$. The material is loaded up to a total strain of one percent. For this numerical experiment, the values of the material parameters are as follows:
@@ 4 Numerical Experiment
[…]
The following figures illustrate the evolution of the internal variables involved in the constitutive model:
Figure 3: Evolution of internal variables (a) $\alpha_{1}$, (b) $r_{1}$, and (c) $k$ over the loading iterations
The evolution of the internal variables is smooth and free from numerical defects. This confirms the stability of the numerical implementation. The backstress $\alpha_{1}$ evolves continuously, reflecting the kinematic hardening, while $r_{1}$ captures the distortion effects. Both attains a saturation value after some point.
Figure 4: Evolution and distortion of the yield surface in the $\sigma_{11}$$\sigma_{12}$ plane.
Figure 4 illustrates the yield surface evolution under monotonic loading. The model captures complex distortional behavior including sharpening in one direction and flattening in the opposite. Such behavior is driven by the evolution of the distortion tensor $r$. The formulation allows independent control over translation, distortion, and rotation in stress space—providing greater flexibility.

### v3

@@ Abstract
This study proposes a modification to the yield condition that addresses the mathematical constraints inherent in the Directional Distortional Hardening models developed by Feigenbaum and Dafalias. The modified model resolves both the mathematical inconsistency found in the ‘complete model’ and the limitations of the ‘r-model’. In the complete model, inconsistency arises between the distortional term in the yield surface and the plastic part of the free energy in the absence of kinematic hardening. Additionally, the ‘r-model’ fails to capture the flattening of the yield surface in the reverse loading direction due to the absence of a fourth-order anisotropic tensor structure in the distortional term. To address these issues, the proposed model introduces a decoupled distortional hardening term in the yield function. This modification enables the simultaneous representation of both flattening and sharpening of the yield surface, and permits isotropic hardening with distortion even without kinematic hardening. A consistent mathematical formulation based on rational thermodynamics and a corresponding numerical algorithm are also developed, establishing a foundation for future experimental investigations and model validation.
@@
Dedicated to Jacob Lubliner
@@ 1 Introduction
Various attempts have been made over the years to mathematically model the yield surface to explain the phenomenon of Directional Distortional Hardening. One of the most notable works is by Baltov and Sawczuk [1] where the author introduced a fourth-order anisotropic tensor in the functional representation of yield surface. This fourth-order anisotropic tensor is the internal variable used to model the distortion. Dafalias et al. [4] proposed a yield surface model, where anisotropy is developed through kinematic and distortional hardening and distortion is represented by a fourth-order tensor $\mathbf{H}$. In this model, the evolution equation of the internal variables was developed using the principles of thermodynamics. Voyiadjis and Foroozesh [21] used the same yield surface as Dafalias et al. [4]. However, the fourth-order distortional tensor $\mathbf{H}$ is defined differently. Voyiadjis and Foroozesh [21] used kinematic considerations to develop the evolution equation for this fourth-order tensor unlike the work of Dafalias et al. [4] where thermodynamic considerations were used in this purpose. Other attempts to model directional distortion have been made by Ortiz and Popov [16], Kurtyka and Życzkowski [11] and François [7]. Ortiz and Popov [16] used a Fourier expansion of the harmonic function to model the yield surface. Kurtyka and Zyczkowski [11] used five hyperspheres with different radius and centers to model the distortion of the yield surface. François [7] included the directional distortion term within the context of thermodynamics for the first time. Unlike the models relying on fourth-order internal variables, his approach introduced distortion by redefining the stress measure itself using a distorted stress formulation governed by second-order ranked internal variables to describe yield surface evolution under non-proportional loading. Feigenbaum and Dafalias [5] developed a ‘complete model’ based on the work of Dafalias et al. [4]. In the ‘complete model’ [5], Feigenbaum incorporated the directionality of the distortion into the yield surface equation. The equations used in this model are derived based on the second law of thermodynamics, specifically the dissipation inequality. This model suggests that energy is released during directional distortion, stored during kinematic and isotropic hardening, and uses evanescent memory type hardening rules. However, the model relies on the back-stress tensor to determine the directionality of distortion, and hence when back-stress equals zero, the model reduces to an isotropic von Mises type model, even if the anisotropy tensor $\mathbf{A}$ is non-zero. And the ‘r-model’ [6] is also developed subsequently by Feigenbaum and Dafalias [6] which introduced an orientational second-order tensor ($\mathbf{r}$) instead of the back-stress ($\boldsymbol{\alpha}$) for the scalar multiplier. The ‘r-model’ replaces the fourth-order distortion tensor $\mathbf{H}$ with the scalar quantity $(1-\mathbf{n^{r}}:\mathbf{r})$, where $\mathbf{n^{r}}$ is the unit radial tensor from the center of the yield surface to the current stress state and $(\mathbf{n^{r}}:\mathbf{r})$ is completely responsible for directional distortion. Therefore, it becomes possible to decouple kinematic hardening from distortional hardening in this model. The mathematical expression for $\mathbf{n}^{r}=\frac{\mathbf{s}-\boldsymbol{\alpha}}{\lVert\mathbf{s}-\boldsymbol{\alpha}\rVert}$, where $\mathbf{s}$ is the deviatoric part of the stress tensor and $\boldsymbol{\alpha}$ is the backstress tensor. Due to the absence of the fourth-order anisotropic term $\mathbf{A}$ in the equation of the yield surface, the flattening of the yield surface in the opposite direction of loading cannot be captured using this model.
Several other notable studies have investigated distortional hardening theory, emphasizing on computational efficiency and simpler internal variable structures. The distortional hardening was captured with the help of second-rank internal variables [19, 17] demonstrating its suitability to closely capture the shape of the yield surfaces from experiments. Shutov et al. [20] developed a finite-strain viscoplasticity model that incorporates distortional hardening, along with isotropic and kinematic hardening, using a second-rank tensor to represent the evolution of anisotropy in the yield surface. Shutov and Ihlemann[19] further extended these ideas by demonstrating that arbitrary convex yield surfaces can be interpolated using evolving second-rank tensors, facilitating smooth and thermodynamically consistent transitions between experimentally observed yield shapes. Later, Marek [13] investigated numerical implementations of distortional hardening models and briefly proposed a modification that is conceptually similar to the one developed in this work, aiming to distort the yield surface in a simplified and computationally efficient manner. Although the distortion mechanism he outlined, particularly the anisotropic term $\mathbf{A}$= $\mathbf{z}\otimes$$\mathbf{z}$, where $\mathbf{z}$ represents a second-rank deviatoric tensor that defines the main axis of the distorted yield surface, was not further pursued in the dissertation, it exhibits similar asymmetry in the yield surface derived in this study.
@@ 2 Framework of Feigenbaum and Dafalias Distortional Hardening Model
$H_{0ijkl}=\frac{3}{2}I_{ijkl}=\frac{3}{2}\left[\frac{1}{2}\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}\right)-\frac{1}{3}\delta_{ij}\delta_{kl}\right],$ (3)
@@
$K_{p}$ is the plastic modulus, which can be determined from the consistency condition. Here, $\frac{\partial f}{\partial\boldsymbol{\sigma}}$ is the gradient of the yield function. The Macaulay brackets $\left\langle\right\rangle$ are used such that, the loading index $\lambda$ only has an effect when the quantity inside the brackets is positive; otherwise, the bracketed term evaluates to zero regardless of $\lambda$’s value. They serve as a mathematical tool to categorize different loading/unloading conditions based on the sign of the loading index $\lambda$, where $\lambda$>>0 for plastic loading, and $\lambda$=0 for neutral loading case.
@@ 2.1 Complete Model and Its Inconsistency
The ‘complete model’ [5] is based on Coleman-Noll procedure-type rational thermodynamics [8] and includes a fourth-order tensor-valued internal variable ($\mathbf{A}$) that describes the evolving directional distortion. Hardening equations for the internal variables, as given in Eqs. (7) to (9), are derived based on sufficient conditions to satisfy the thermodynamic requirements of positive energy dissipation. Evolution equtions of the internal variables are all fading memory-type equations.
@@
The isotropic, kinematic, and distortional parts of the free energy are positive definite, and their evolution equations asymptotically approach a saturation limit. In addition, if the internal variable of a hardening term is zero, the corresponding free energy also becomes zero. For example, if $\boldsymbol{\alpha}=\boldsymbol{0}$, $\psi_{p}^{\text{kin}}=\frac{1}{2a_{1}}\boldsymbol{\alpha}:\boldsymbol{\alpha}=0$, and Eq. (12)reduces to Eq. (13):
@@
[…]
@@
Because of this simplification, it is possible to model distortion and translation of the yield surface separately and simultaneously, which was a major issue in the complete model. The mathematical form of the yield surface in Eq. (15) is motivated by the experiments of Ishikawa [9], which suggest that the yield surface starts as a ‘von Mises’ hyper-circle, and the scalar term $(\mathbf{n^{r}}:\mathbf{r})$, obtained from double contraction of two second rank tensor, sharpens the yield surface in the direction of loading. However, the flattening in the opposite side of the yield surface, as observed in the experiments [22, 18, 14, 15], cannot be modeled using this ‘r-model’. A yield surface model with a fourth-order anisotropic tensor could provide an effective means to capture this flattening phenomenon.
@@ 3.2 Hardening Rules based on the Proposed Model
Based on the proposed form of the yield surface in Eq. (17), the evolution equations for the model, along with the relevant expressions, are derived below in accordance with the second law of thermodynamics.
@@ 3.2.1 Expression of $\frac{\partial f}{\partial\boldsymbol{\sigma}}$ and Flow Rule
$\frac{\partial f}{\partial\boldsymbol{\sigma}}=||\mathbf{s}-\boldsymbol{\alpha}||\{3\mathbf{n}^{r}+2(\mathbf{n}^{r}:\mathbf{r})(\mathbf{r}\otimes\mathbf{r}):\mathbf{n}^{r}+[\mathbf{n}^{r}:(\mathbf{r}\otimes\mathbf{r})]:\mathbf{n}^{r}(\mathbf{r}-(\mathbf{n}^{r}:\mathbf{r})\mathbf{n}^{r})\}.$ (18)
@@ 3.2.2 Derivation of Hardening Rules
$\psi_{p}^{\text{iso}}=\sum_{i=1}^{l}\frac{\kappa_{i}}{2\rho}k_{c}^{2},$ (21)
$\psi_{p}^{\text{kin}}=\sum_{i=1}^{l}\frac{a_{i}}{2\rho}\boldsymbol{{\alpha_{c}}}:\boldsymbol{{\alpha_{c}}},$ (22)
$\psi_{p}^{\text{dis}}=\sum_{i=1}^{l}\frac{A_{i}}{2\rho}\mathbf{r_{c}}:\mathbf{r_{c}}.$ (23)
To keep the calculation simple and the number of parameters to the minimum, only the first term is considered in Eqs. (21)–(23), i.e., $l=1$, and no superposition is considered. The final expressions are given in Eqs. (24)–(26):
@@
$\dot{\boldsymbol{\varepsilon}_{p}}-\frac{1}{a_{1}}\dot{\boldsymbol{\alpha}}=\lambda\left\|\frac{\partial f}{\partial\boldsymbol{\sigma}}\right\|a_{2}\boldsymbol{\alpha}\quad\implies\quad\dot{\boldsymbol{\alpha}}=\lambda\left\|\frac{\partial f}{\partial\boldsymbol{\sigma}}\right\|a_{1}\left(\mathbf{n}-a_{2}\boldsymbol{\alpha}\right),$ (35)
$\lambda\,||\mathbf{s}-\boldsymbol{\alpha}||^{2}\Bigl[(\mathbf{n^{r}}:\mathbf{r})\,\mathbf{n^{r}}:(\mathbf{r}\otimes\mathbf{r}):\mathbf{n^{r}}\Bigr]+\frac{1}{A_{1}}\,\mathbf{r}:\boldsymbol{\dot{r}}=-\frac{3}{2}\lambda\,||\mathbf{s}-\boldsymbol{\alpha}||^{2}\,A_{2}\,(\mathbf{r}:\mathbf{r}),\\$
@@
The rate equations obtained are of the evanescent memory type, i.e., these are the Armstrong-Frederick model type equations. This means the internal variables have an upper limit on their evolution and are path-dependent. For Eq. (36) to satisfy Eq. (33), the necessary condition is given in Eq. (40):
@@ 3.2.3 Plastic Modulus
$\frac{\partial f(\boldsymbol{\sigma},\boldsymbol{\alpha},\mathbf{r},k)}{\partial\boldsymbol{\alpha}}=-\frac{\partial f(\boldsymbol{\sigma},\boldsymbol{\alpha},\mathbf{r},k)}{\partial\boldsymbol{\sigma}}.$ (43)
@@ 3.2.4 Constraints on Material Parameters
Eq. (50): puts a constraint on the upper bound of $A_{2}$. Another way to find and cross-check the limit is by substituting Eqs. (4749) into the expression of the gradient of the yield surface at the limiting condition. Substituting the limiting values of the internal variables obtained in Eqs. (4749) into the expression of the gradient of the yield surface given in Eq. (18), the limiting value of the gradient is obtained and given in Eq. (51):
@@
Remarkably, the thermodynamic approach and the convexity requirement of the yield surface both agree upon the restriction put on the material constant. In addition, the lower bound for $A_{2}$ can be calculated from the requirement of positive definiteness of the equivalent stress. This condition, evaluated at the limiting case of $\mathbf{n}^{r}$ and $\boldsymbol{\alpha}$, leads to the expression given in Eq. (54).
@@
Which puts a lower bound on $A_{2}$. However, if $\mathbf{n}^{r}$ and $\mathbf{n}^{rl}$ are in the same direction (i.e., $\mathbf{n}^{r}:\mathbf{n}^{r^{l}}=+1$), then Eqs. (50) and (53) are recovered.
@@
4 Numerical Implementation
The goal of this section is to show that the theoretical model can be implemented numerically using the algorithm described in the Appendix A. To show this the material is loaded uniaxially with a strain increment of $5\times 10^{-5}$. The material is loaded up to a total strain of one percent. For this numerical experiment, the values of the material parameters are as follows:
@@ 4 Numerical Implementation
$\left.\begin{aligned} \dot{k}&=\lambda k\kappa_{1}(1-\kappa_{2}k),\quad\dot{\boldsymbol{\alpha}}=\lambda\left\|\frac{\partial f}{\partial\boldsymbol{\sigma}}\right\|a_{1}\left(\mathbf{n}-a_{2}\boldsymbol{\alpha}\right),\\ \dot{\mathbf{r}}&=-\lambda\,A_{1}\,\|\mathbf{s}-\boldsymbol{\alpha}\|^{2}\left[(\mathbf{n}^{r}:\mathbf{r})^{2}\,\mathbf{n}^{r}+\frac{3}{2}A_{2}\,\mathbf{r}\right].\end{aligned}\right\}$ (60)
The following figures in Fig. 3 illustrate the evolution of the internal variables involved in the constitutive model:
Figure 3: Evolution of internal variables (a) $k$, (b) $\alpha_{1}$, and (c) $r_{1}$ over the loading iterations
The evolution of the internal variables is smooth and free from numerical defects. This confirms the stability of the numerical implementation. The backstress $\alpha_{1}$ evolves continuously, reflecting the kinematic hardening, while $r_{1}$ captures the distortion effects.
Figure 4: Evolution and distortion of the modeled Yield Surface (Y.S.) in the $\sigma_{11}$$\sigma_{12}$ plane.
Fig. 4 illustrates the yield surface evolution under monotonic loading. The model captures complex distortional behavior including sharpening in one direction and flattening in the opposite. Such behavior is driven by the evolution of the distortion tensor $\boldsymbol{r}$. The formulation allows independent control over translation, distortion, and rotation in the stress space, providing greater flexibility.